OpenStudy (anonymous):

TRUE/FALSE The equation cosx+sinx= Radical 3 +1 over 2 is an identity.

OpenStudy (anonymous):

\[cosx+sinx = \frac{ \sqrt{3}+1 }{ 2 }\]

OpenStudy (anonymous):

@hartnn @marissalovescats do you know this

OpenStudy (anonymous):

It's false....

hartnn (hartnn):

yeah, the only identity relating sin x and cos x only is \(\large \sin^2 x+\cos^2x =1\)